Circular No. 7116

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

4U 0115+63 R. B. Wilson and B. A. Harmon, Marshall Space Flight Center, NASA;

and M. H. Finger, Universities Space Research Association, report for the Compton GRO BATSE team: “An outburst is currently being detected from the Be/x-ray pulsar 4U 0115+63. Pulsations from the source were first detected on Feb. 22 at a frequency of 0.27666170(75) Hz, assuming the binary-orbit ephemeris of Bildsten et al. (1997, Ap.J. Suppl. 113, 367). The rms average pulsed fluxes in the energy range 20–50 keV on Feb. 22 and 23 were 0.132 ± 0.030 and 0.177 ± 0.024 keV cm^{−2} s^{−1}, respectively. The source is detected by earth occultation after Feb. 18, reaching an average intensity between Feb. 20 and 23 of 0.31 ± 0.08 keV cm^{−2} s^{−1} in the band 20–50 keV.”

Corrigenda. On IAUC 3510, 5990, 5999, 6266, 6272, 6450, and 6451, this object was erroneously given as “4U 0115+634”.

1ES 1426+428 G. Ghisellini, G. Tagliaferri, Osservatorio Astronomico di Brera, Mi-

lan; and P. Giommi, BeppoSAX, Rome, report: “The BL Lac object 1ES 1426+428 was observed by BeppoSAX starting on Feb. 8.865 UT. Preliminary results indicate a hard x-ray spectrum, with energy spectral index ∼ 0.85 in the band 0.1–10 keV. The spectrum at higher energies, up to 100 keV, shows an excess with respect to the extrapolation from the band 2–10 keV. The 2–10-keV flux is ∼ 2 × 10^{−11} erg cm^{−2} s^{−1}. Examination of the All Sky Monitor XTE light curve shows that the source remained in a very active x-ray state also after the BeppoSAX observation. Since the entire x-ray spectrum of this source is believed to be due to the synchrotron process, we foresee an intense TeV emission. Observations in the TeV and optical bands are strongly encouraged.”

SUPERNOVA 1999aa IN NGC 2595 Magnitude estimates: Feb. 15.855 UT, 14.9 (D. Hanžl, Brno, Czech

Republic; CCD); 17.911, 15.0 (K. Hornoch, Lelekovice, Czech Republic; visual); 18.897, 14.8 (Hornoch); 22.698, 15.2 (S. Yoshida and K. Kadota, Ageo City, Japan; CCD); 23.810, 14.8 (Hornoch).

1999 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green